# VisualQuality-R1: Reasoning-Induced Image Quality Assessment via Reinforcement Learning to Rank

### v1

@@ Abstract
DeepSeek-R1 has demonstrated remarkable effectiveness in incentivizing reasoning and generalization capabilities of large language models (LLMs) through reinforcement learning. Nevertheless, the potential of reasoning-induced computational modeling has not been thoroughly explored in the context of image quality assessment (IQA), a task critically dependent on visual reasoning. In this paper, we introduce VisualQuality-R1, a reasoning-induced no-reference IQA (NR-IQA) model, and we train it with reinforcement learning to rank, a learning algorithm tailored to the intrinsically relative nature of visual quality. Specifically, for a pair of images, we employ group relative policy optimization to generate multiple quality scores for each image. These estimates are then used to compute comparative probabilities of one image having higher quality than the other under the Thurstone model. Rewards for each quality estimate are defined using continuous fidelity measures rather than discretized binary labels. Extensive experiments show that the proposed VisualQuality-R1 consistently outperforms discriminative deep learning-based NR-IQA models as well as a recent reasoning-induced quality regression method. Moreover, VisualQuality-R1 is capable of generating contextually rich, human-aligned quality descriptions, and supports multi-dataset training without requiring perceptual scale realignment. These features make VisualQuality-R1 especially well-suited for reliably measuring progress in a wide range of image processing tasks like super-resolution and image generation.
@@ 1 Introduction
The rapid advancement of vision-language models (VLMs) offers promising avenues for enhancing NR-IQA generalization by contextualizing it into broader vision tasks wu2024comprehensive . VLMs can effectively integrate multi-modal information, enabling understanding of both low-level image distortions, such as noise and blur, and high-level perceptual attributes, such as aesthetics and content semantics. This multi-modal semantic contextualization enables VLMs to articulate nuanced quality descriptions with stronger generalization.
@@
Reinforcement learning (RL) has recently emerged as a powerful alternative, enhancing the reasoning capabilities of LLMs, while aligning their responses with human preferences rafailov2023direct ; guo2025deepseek . In particular, DeepSeek-R1 guo2025deepseek demonstrates the effectiveness of RL in promoting generalization by encouraging automated exploration of plausible reasoning paths and employing rule-based rewards to prevent reward hacking skalse2022defining . However, a direct adaptation of RL techniques to NR-IQA, as exemplified by the recent Q-Insight model li2025qinsight , has been limited by its reliance on dataset-specific reward design and additional distortion-type classification. These constraints stem from its treatment of visual quality as an absolute perceptual quantity, thereby framing NR-IQA simplistically as a regression task.
In this paper, we introduce VisualQuality-R1, a reasoning-induced NR-IQA model, and we train it via reinforcement learning to rank (RL2R), a learning algorithm explicitly designed to capture the inherently relative nature of visual quality. Specifically, we employ group relative policy optimization (GRPO) shao2024deepseekmath to derive multiple quality scores for each image in a pair. We then compute comparative probabilities between images using the Thurstone model thurstone1927law by assessing the difference between the mean quality score of one image and individual quality scores of another, normalized by their added sample variance. Unlike previous methods, we define the reward function using the continuous fidelity measure tsai2007frank , which provides precise guidance to facilitate quality ranking. Extensive experiments confirm that VisualQuality-R1 effectively assesses visual quality across a diverse range of distortion scenarios, outperforming discriminative deep learning-based NR-IQA models as well as a recent reasoning-induced quality regression method. Moreover, VisualQuality-R1 generates contextually rich, human-aligned quality descriptions (see Fig. 1), which can be leveraged to provide targeted feedback for downstream image processing algorithms, and support fine-grained quality control in digital photography pipelines. Additionally, we demonstrate that VisualQuality-R1 remains effective across multi-dataset training scenarios without requiring perceptual scale realignment. As a result, VisualQuality-R1 represents a substantial advancement in NR-IQA, and establishes new benchmarks for reasoning and generalization performance.
@@ 2 Related Work
NR-IQA models primarily employed regression-based approaches, wherein image quality was treated as an absolute perceptual quantity directly estimated from extracted “quality-aware” features. Initially, features were handcrafted based on natural scene statistics mittal2012no ; mittal2012making , degradation-specific characteristics wang2000blind ; wang2003local ; liu2009no , and perceptual models inspired by the human visual system wang2011reduced . However, these NR-IQA methods were limited by the representational capacity of handcrafted features. Later, deep learning-based regression models emerged as the dominant paradigm, using end-to-end trainable neural networks to directly predict quality scores (or, in some cases, quality distributions) kang2014convolutional ; ma2017end ; bosse2017deep ; talebi2018nima ; ke2021musiq ; yang2022maniqa . These models typically utilize standard regression losses such as the mean squared error and mean absolute error, or statistical distances such as the earth mover’s distance talebi2018nima and Kullback–Leibler (KL) divergence you2025teaching . Regression-based models often struggle with generalization issues, and require labor-intensive perceptual scale realignment mikhailiuk2021consolidated when training on multiple IQA datasets.
@@
To address these shortcomings, ranking-based NR-IQA models were introduced, modeling visual quality as an intrinsically relative perceptual quantity. Gao et al. gao2015learning pioneered the concept of quality ranking in NR-IQA, although their initial implementation relied on predefined anchor images and was not end-to-end optimized. Ma et al. ma2017dipiq adapted RankNet burges2005learning to NR-IQA by training (though not fully end-to-end) on quality-discriminable image pairs. Their subsequent work established the first end-to-end ranking-based NR-IQA method grounded in the Thurstone model thurstone1927law . Nonetheless, their approach suffers from scaling ambiguity during variance estimation. Zhang et al. zhang2021uncertainty incorporated a hinge loss to regularize variance estimation, yet the scaling ambiguity persisted. Their study demonstrated the superiority of the fidelity loss tsai2007frank over the conventional cross-entropy loss in ranking-based NR-IQA. Subsequent research has adopted a simpler approach by fixing the variance parameter to one (corresponding to the Thurstone Case V model), facilitating active fine-tuning of NR-IQA models on challenging examples wang2021troubleshooting ; wang2021active and allowing continual adaptation to novel distortion scenarios zhang2022continual . Other losses that enable quality ranking include the margin ranking loss liu2017rankiqa , differentiable approximations of Spearman’s rank correlation coefficient (SRCC) blondel2020fast , Pearson linear correlation coefficient (PLCC) xu2024boosting , and statistical distances between permutation probabilities cao2007learning ; stern1990models ; irurozki2019mallows .
@@
The integration of VLMs into NR-IQA has recently gained traction, particularly due to their proficiency in capturing semantic context through multi-modal representation learning. Early attempts include multitask adaptation of CLIP zhang2023blind , as well as SFT-based methods like Q-Align wu2024qalign , Compare2Score zhu2024adaptive , DepictQA you2024depicting , and DeQA-Score you2025teaching , which trained VLMs to generate either quality scores, distributions, or descriptions. Closest to ours, Q-Insight li2025qinsight explored reasoning-induced quality regression through RL. However, Q-Insight struggles with dataset-specific reward calibration, generalization to novel distortion scenarios, and the added complexity of auxiliary distortion-type classification. In contrast, our VisualQuality-R1 redefines the use of VLMs in NR-IQA by shifting from absolute regression to relative ranking, leading to enhanced generalization across distortion scenarios with better quality justifications.
Figure 2: System diagram of the proposed VisualQuality-R1 trained via RL2R. Given an image pair $(x_{i},x_{j})$ with a shared text prompt $c$, VisualQuality-R1 generates $K$ responses. Following GRPO shao2024deepseekmath , each response includes a detailed reasoning process and a predicted quality score. To assess relative visual quality, we calculate the asymmetric comparative probability that image $x_{i}$ is perceived better than $x_{j}$ under the Thurstone model thurstone1927law . This involves subtracting the mean predicted score of $x_{j}$ from the $k$-th score of $x_{i}$, standardized by their added sample variance. A fidelity reward is derived from human preference, providing continuous supervisory signals for policy optimization.
@@ 3 Reasoning-Induced NR-IQA
To harness both the powerful reasoning-inducing capabilities of RL and the intrinsically relative nature of visual quality, we propose an NR-IQA model—VisualQuality-R1—and an RL2R method of training it that seamlessly integrates the Thurstone model within GRPO. Fig. 2 shows the system diagram of VisualQuality-R1 trained via RL2R.
@@ 3.1 VisualQuality-R1 via RL2R
More specifically, for a training batch of images $\{x_{1},x_{2},\ldots,x_{B}\}$, where $B$ is the minibatch size, we apply GRPO to generate $K$ quality predictions for $x_{i}$, $q(x_{i})=[q_{1}(x_{i}),q_{2}(x_{i}),\ldots,q_{K}(x_{i})]^{\intercal}$. This output naturally encodes predictive uncertainty, which is crucial for making reliable relative quality ranking. Under the Thurstone model thurstone1927law , the visual quality of an image is assumed to follow a Gaussian distribution. Thus, we are able to compute the asymmetric comparative probability for each of the $B\times(B-1)$ ordered image pairs by subtracting the mean quality score of $x_{j}$ from the $k$-th quality score of $x_{i}$, standardized by their added sample variance:
@@
Rate the overall image visual quality. The rating should be a float between 1 and 5, with 1 representing very poor quality and 5 representing excellent quality. First output the thinking process in <think> </think> tags and then output the final answer with only one score in <answer> </answer> tags.
An important aspect of our RL2R algorithm is that we define the reward function $r_{k}(x_{i})$ for each quality estimate $q_{k}(x_{i})$ as the fidelity measure tsai2007frank —a continuous analogue of the discretized binary reward guo2025deepseek ; li2025qinsight , averaged across all $B-1$ image pairs:
@@
This continuous reward feedback provides precise guidance during RL2R by capturing subtle distinctions in quality ranking, thus boosting generalization across diverse distortion scenarios. We collect $K$ fidelity rewards for $x_{i}$ into the vector $r(x_{i})=[r_{1}(x_{i}),r_{2}(x_{i}),\ldots,r_{K}(x_{i})]^{\intercal}$, and compute the relative advantage $a_{k}(x_{i})$ by standardizing rewards within group:
@@
We conclude this section by highlighting the key strengths of the resulting VisualQuality-R1. First, VisualQuality-R1 inherits all the advantages of ranking-based NR-IQA models, enabling effective multi-dataset training, active fine-tuning, and continual model adaptation without requiring perceptual scale realignment mikhailiuk2021consolidated , a feature notably absent in regression-based NR-IQA approaches. Second, trained via RL2R, VisualQuality-R1 mitigates the scalability and overfitting issues inherent in SFT-based models. Third, VisualQuality-R1 promises to both improve model generalizability and furnish contextually rich textual justifications alongside numerical quality scores, thereby boosting its practical relevance in real-world IQA applications.
@@ 4.1 Experimental Setups
We fine-tune Qwen2.5-VL-7B bai2025qwen2 as the backbone for VisualQuality-R1 using GRPO shao2024deepseekmath , without freezing any layers. The AdamW optimizer loshchilov2017decoupled is employed with an initial learning rate of $1\times 10^{-6}$ and a linear decay schedule. For GRPO, we generate six candidate responses per prompt (i.e., $K=6$) and set the balance coefficient $\beta$ to $0.04$. Training runs on $16$ NVIDIA A100 GPUs with a minibatch size of eight per GPU, taking approximately five hours for a total of $10$ epochs.
@@ 4.2 Main Results
We first train NR-IQA models exclusively on the synthetic KADID-10K lin2019kadid training set ($6:2:2$ split while ensuring content independence) and tested in a zero-shot setting across eight datasets with distortions arising from digital imaging and (post-)processing stages: BID ciancio2010no , CLIVE ghadiyaram2015massive , KonIQ-10k hosu2020koniq , SPAQ fang2020perceptual , Liu13 (deblurring) liu2013no , SRIQA-Bench (super-resolution) chen2025toward , Min19 (dehazing) min2019quality , and AGIQA-3K (image generation) li2023agiqa .
@@
Our RL2R approach enables multi-dataset training without the need for perceptual scale realignment. To exploit this, we train VisualQuality-R1 on a combination of KADID-10K lin2019kadid and SPAQ fang2020perceptual (again $6:2:2$ split while ensuring content independence). As shown in Table 2, VisualQuality-R1 yields consistent performance gains. Despite a minor dip in the image generation scenario, the average SRCC/PLCC rises from $0.777/0.814$ to $0.791/0.831$. In stark contrast, Q-Insight li2025qinsight fails to benefit from multi-dataset training due to its inability to address perceptual scale variations [^2] To enable multi-dataset training of Q-Insight, we linearly rescale MOSs from different IQA datasets to $[1,5]$, and apply a dataset-agnostic threshold to compute binary rewards.: KADID-10K uses ratings from $1$ to $5$, while SPAQ spans 00 to $100$.
@@
# of Generated Responses Imaging-Related Distortion Processing-Related Distortion Avg

### v2

@@ Abstract
DeepSeek-R1 has demonstrated remarkable effectiveness in incentivizing reasoning and generalization capabilities of large language models (LLMs) through reinforcement learning. Nevertheless, the potential of reasoning-induced computation has not been thoroughly explored in the context of image quality assessment (IQA), a task depending critically on visual reasoning. In this paper, we introduce VisualQuality-R1, a reasoning-induced no-reference IQA (NR-IQA) model, and we train it with reinforcement learning to rank, a learning algorithm tailored to the intrinsically relative nature of visual quality. Specifically, for a pair of images, we employ group relative policy optimization to generate multiple quality scores for each image. These estimates are used to compute comparative probabilities of one image having higher quality than the other under the Thurstone model. Rewards for each quality estimate are defined using continuous fidelity measures rather than discretized binary labels. Extensive experiments show that the proposed VisualQuality-R1 consistently outperforms discriminative deep learning-based NR-IQA models as well as a recent reasoning-induced quality regression method. Moreover, VisualQuality-R1 is capable of generating contextually rich, human-aligned quality descriptions, and supports multi-dataset training without requiring perceptual scale realignment. These features make VisualQuality-R1 especially well-suited for reliably measuring progress in a wide range of image processing tasks like super-resolution and image generation.
@@ 1 Introduction
The rapid advancement of vision-language models (VLMs) offers promising avenues for enhancing NR-IQA generalization by contextualizing it into broader vision tasks wu2024comprehensive . VLMs can effectively integrate multi-modal information, enabling understanding of both low-level image distortions (e.g., noise and blur) and high-level perceptual attributes (e.g., aesthetics and content semantics). This multi-modal semantic contextualization allows VLMs to articulate nuanced quality descriptions with stronger generalization.
@@
Reinforcement learning (RL) has recently emerged as a powerful alternative, enhancing the reasoning capabilities of LLMs, while aligning their responses with human preferences rafailov2023direct ; guo2025deepseek . In particular, DeepSeek-R1 guo2025deepseek demonstrates the effectiveness of RL in promoting generalization by encouraging automated exploration of plausible reasoning paths and employing rule-based rewards to prevent reward hacking skalse2022defining . However, a direct adaptation of RL techniques to NR-IQA, as exemplified by the recent Q-Insight model li2025qinsight , has been limited by its reliance on dataset-specific reward design and additional distortion-type classification. These constraints stem from its treatment of visual quality as an absolute perceptual quantity, thereby framing NR-IQA naïvely as a regression task.
In this paper, we introduce VisualQuality-R1, a reasoning-induced NR-IQA model, and we train it via reinforcement learning to rank (RL2R), a learning algorithm explicitly designed to capture the inherently relative nature of visual quality. Specifically, we employ group relative policy optimization (GRPO) shao2024deepseekmath to derive multiple quality scores for each image in a pair. We then compute comparative probabilities between images using the Thurstone model thurstone1927law by assessing the difference between the mean quality score of one image and individual quality scores of another, normalized by their sample variances. Unlike previous methods, we define the reward function using the continuous fidelity measure tsai2007frank , which provides precise guidance to facilitate quality ranking. Extensive experiments confirm that VisualQuality-R1 effectively assesses visual quality across a diverse range of distortion scenarios, outperforming discriminative deep learning-based NR-IQA models as well as a recent reasoning-induced quality regression method li2025qinsight . Moreover, VisualQuality-R1 generates contextually rich, human-aligned quality descriptions (see Fig. 1), which can be leveraged to provide targeted feedback for downstream image processing algorithms, and support fine-grained quality control in digital photography pipelines. Additionally, we demonstrate that VisualQuality-R1 remains effective across multi-dataset training scenarios without requiring perceptual scale realignment.
@@ 2 Related Work
NR-IQA models primarily employed regression-based approaches, wherein image quality was treated as an absolute perceptual quantity directly estimated from extracted “quality-aware” features. Initially, features were handcrafted based on natural scene statistics mittal2012no ; mittal2012making , degradation-specific characteristics wang2000blind ; wang2003local ; liu2009no , and perceptual models inspired by the human visual system wang2011reduced . Nonetheless, these methods were limited by the representational capacity of handcrafted features. Later, deep learning-based regression models emerged as the dominant paradigm, using end-to-end trainable neural networks to directly predict quality scores (or, in some cases, quality distributions) kang2014convolutional ; ma2017end ; bosse2017deep ; talebi2018nima ; yang2022maniqa ; wu2023assessor360 . These models typically utilize standard regression losses such as the mean squared error and mean absolute error, or statistical distances such as the earth mover’s distance talebi2018nima and Kullback–Leibler (KL) divergence you2025teaching . Regression-based models often struggle with generalization issues, and require labor-intensive perceptual scale realignment mikhailiuk2021consolidated when training on multiple IQA datasets.
@@
To address these shortcomings, ranking-based NR-IQA models were introduced, modeling visual quality as an intrinsically relative perceptual quantity. Gao et al. gao2015learning pioneered the concept of quality ranking in NR-IQA, although their initial implementation relied on predefined anchor images and was not end-to-end optimized. Ma et al. ma2017dipiq adapted RankNet burges2005learning to NR-IQA by training (though not fully end-to-end) on quality-discriminable image pairs. Their subsequent work established the first end-to-end ranking-based NR-IQA method grounded in the Thurstone model thurstone1927law . Nevertheless, their approach suffers from scaling ambiguity during variance estimation. Zhang et al. zhang2021uncertainty incorporated a hinge loss to regularize variance estimation, yet the scaling ambiguity persisted. Their study also demonstrated the superiority of the fidelity loss tsai2007frank over the conventional cross-entropy loss in ranking-based NR-IQA. Subsequent research has adopted a simpler approach by fixing the variance parameter to one (corresponding to the Thurstone Case V model), facilitating active fine-tuning of NR-IQA models on challenging examples wang2021troubleshooting ; wang2021active and allowing for continual adaptation to novel distortion scenarios zhang2022continual . Other losses that enable quality ranking include the margin ranking loss liu2017rankiqa , differentiable approximations of Spearman’s rank correlation coefficient (SRCC) blondel2020fast , Pearson linear correlation coefficient (PLCC) xu2024boosting , and statistical distances between permutation probabilities cao2007learning ; stern1990models ; irurozki2019mallows .
@@
The integration of VLMs into NR-IQA has recently gained traction, particularly due to their proficiency in capturing contextual semantics through multi-modal representation learning. Early attempts include multitask adaptation of CLIP zhang2023blind , as well as SFT-based methods like Q-Align wu2024qalign , Compare2Score zhu2024adaptive , DepictQA you2024depicting , and DeQA-Score you2025teaching , which trained VLMs to generate either quality scores, distributions, or descriptions. Closest to ours, Q-Insight li2025qinsight explored reasoning-induced quality regression through RL. However, Q-Insight struggles with the dataset-specific reward calibration, added complexity of auxiliary distortion-type classification, and generalization to novel distortion scenarios. In contrast, our VisualQuality-R1 redefines the use of VLMs in NR-IQA by shifting from absolute regression to relative ranking, leading to enhanced generalization across distortion scenarios with better quality justifications.
Figure 2: System diagram of the proposed VisualQuality-R1 trained via RL2R. Given an image pair $(x_{i},x_{j})$ with a shared text prompt $c$, VisualQuality-R1 generates $K$ responses. Following GRPO shao2024deepseekmath , each response includes a detailed reasoning process and a predicted quality score. To assess relative visual quality, we calculate the asymmetric comparative probability that image $x_{i}$ is perceived better than $x_{j}$ under the Thurstone model thurstone1927law . This involves subtracting the mean predicted score of $x_{j}$ from the $k$-th score of $x_{i}$, standardized by their sample variances. A fidelity reward is derived from human preference, providing continuous supervisory signals for policy optimization.
@@ 3 Reasoning-Induced NR-IQA
To harness both the powerful reasoning-inducing capabilities of RL and the intrinsically relative nature of visual quality, we propose an NR-IQA model—VisualQuality-R1—and an RL2R method of training it that seamlessly integrates the Thurstone model within GRPO. Fig. 2 shows the system diagram of VisualQuality-R1.
@@ 3.1 VisualQuality-R1 via RL2R
More specifically, for a training batch of images $\{x_{1},x_{2},\ldots,x_{B}\}$, where $B$ is the minibatch size, we apply GRPO to generate $K$ quality predictions for $x_{i}$, $q(x_{i})=[q_{1}(x_{i}),q_{2}(x_{i}),\ldots,q_{K}(x_{i})]^{\intercal}$. This output naturally encodes predictive uncertainty, which is crucial for making reliable relative quality ranking. Under the Thurstone model thurstone1927law , the visual quality of an image is assumed to follow a Gaussian distribution. Thus, we compute the asymmetric comparative probability for each of the $B\times(B-1)$ ordered image pairs by subtracting the mean quality score of $x_{j}$ from the $k$-th quality score of $x_{i}$, standardized by their sample variances:
@@
What is your overall rating on the quality of this picture? The rating should be a float between 1 and 5, rounded to two decimal places, with 1 representing very poor quality and 5 representing excellent quality. First output the thinking process in <think> </think> tags and then output the final answer with only one score in <answer> </answer> tags.
An important aspect of our RL2R algorithm is that we define the reward function $r_{k}(x_{i})$ for each quality estimate $q_{k}(x_{i})$ as the fidelity measure tsai2007frank —a continuous analogue of the discretized binary reward guo2025deepseek ; li2025qinsight , averaged across all B1B-1 image pairs:
@@
This continuous reward feedback provides precise guidance during RL2R by capturing subtle variations in quality ranking, thus improving generalization across diverse distortion scenarios. We collect $K$ fidelity rewards for $x_{i}$ into the vector $r(x_{i})=[r_{1}(x_{i}),r_{2}(x_{i}),\ldots,r_{K}(x_{i})]^{\intercal}$, and compute the relative advantage $a_{k}(x_{i})$ by standardizing rewards within group:
@@
We conclude this section by highlighting the key strengths of our VisualQuality-R1. First, VisualQuality-R1 inherits all the advantages of ranking-based NR-IQA models, enabling effective multi-dataset training, active fine-tuning, and continual model adaptation without requiring perceptual scale realignment mikhailiuk2021consolidated , a feature notably absent in regression-based NR-IQA approaches. Second, trained via RL2R, VisualQuality-R1 mitigates the scalability and overfitting issues inherent in SFT-based models. Third, VisualQuality-R1 promises to both improve model generalizability and furnish contextually rich textual justifications alongside numerical quality scores, thereby boosting its practical relevance in real-world IQA applications.
@@ 4.1 Experimental Setups
We fine-tune Qwen2.5-VL-7B bai2025qwen2 as the backbone for VisualQuality-R1 using GRPO shao2024deepseekmath . The AdamW optimizer loshchilov2017decoupled is employed with an initial learning rate of $1\times 10^{-6}$ and a linear decay schedule. For GRPO, we generate six candidate responses per prompt (i.e., $K=6$) and set the balance coefficient $\beta$ to $0.04$. Training runs on $16$ NVIDIA A100 GPUs with a minibatch size of eight per GPU, taking approximately five hours for a total of $10$ epochs.
@@ 4.2 Main Results
We first train NR-IQA models on the synthetic KADID-10K lin2019kadid training set ($6:2:2$ split while ensuring content independence) and test in a zero-shot setting across eight datasets with distortions arising from digital imaging and (post-)processing stages: BID ciancio2010no , CLIVE ghadiyaram2015massive , KonIQ-10k hosu2020koniq , SPAQ fang2020perceptual , Liu13 (deblurring) liu2013no , SRIQA-Bench (super-resolution) chen2025toward , Min19 (dehazing) min2019quality , and AGIQA-3K (image generation) li2023agiqa .
@@
Our RL2R approach enables multi-dataset training without the need for perceptual scale realignment. To exploit this, we train VisualQuality-R1 on a combination of KADID-10K lin2019kadid and SPAQ fang2020perceptual (again $6:2:2$ split while ensuring content independence). As shown in Table 2, VisualQuality-R1 yields consistent performance gains. Despite a minor dip in the image generation scenario, the average SRCC/PLCC rises from $0.777/0.814$ to $0.791/0.831$. In stark contrast, Q-Insight li2025qinsight fails to benefit from multi-dataset training due to its inability to address perceptual scale variations [^2] To enable multi-dataset training of Q-Insight, we linearly rescale MOSs from different IQA datasets to $[1,5]$, and apply a dataset-agnostic threshold to compute binary rewards.: KADID-10K uses ratings from $1$ to $5$, while SPAQ spans 0 to $100$.
@@
#Generated Responses Imaging-Related Distortion Processing-Related Distortion Avg